# Benchmarking Differentially Private Tabular Data Synthesis

### v1

@@ 3.1. Existing Algorithms
More recent machine learning approaches, including GEM (Liu et al., 2021) and DP-MERF (Harder et al., 2021), represent generative network-based advancements. GEM combines generative networks with adaptive marginal selection mechanisms, while DP-MERF employs random Fourier feature loss to train generative networks. Besides, TabDDPM (Kotelnikov et al., 2023) leverages diffusion models’ representational power to fit target data directly. While TabDDPM was not originally designed for DP, it achieves state-of-the-art performance among non-DP methods and can be adapted for DP synthesis using DP-SGD (Du and Li, 2024). Consequently, we include TabDDPM in our analysis.
Our work focuses on recently proposed methods that have not been well compared previously and those representing the current state-of-the-art, as summarized in Table 1. Here, we must emphasize that specifically, we don’t consider some LLM-based methods (Tran and Xiong, 2024; Sablayrolles et al., 2023). That is because LLMs are trained on extensive public datasets, which will introduce evaluation bias and affect our further comparison of algorithm modules.
3.2. Existing Benchmark Works
In addition to the proposed algorithms, there are several benchmark studies (Du and Li, 2024; Hu et al., 2023; Yang et al., 2024; Fan, 2020; Tao et al., 2022) that investigate the problem of DP data synthesis. However, previous works all have some weaknesses, summarized as follows.
  • Lack of unified comparison setting. Du et al.  (Du and Li, 2024) have made an experimental evaluation of current works but ignore the importance of a unified setting (e.g., preprocessing), which may significantly influence the comparison fairness.
  • Not include recent advanced works. Du et al.  (Du and Li, 2024) and Tao et al. (Tao et al., 2022) focus on utility evaluation, but they both lack investigation for some advanced methods, such as RAP++ and Private-GSD. Fan et al.’s work  (Fan, 2020) only focuses on GAN-based methods.
  • Lack of deep analysis. Yang et al. (Yang et al., 2024) provide a survey of many methods and further investigate distributed data synthesis. However, they do not delve deep into these algorithms’ working principles. Du et al.’s work and Tao et al.’s work also have a similar weakness in lacking deep algorithm analysis.
  • Lack of comprehensive experiments. Hu et al. (Hu et al., 2023) provide analysis for a wide range of DP synthesis algorithms, not only about tabular data synthesis but also trajectory data and graph data. This work, though trying to make an in-depth analysis, lacks comprehensive experiments.
Except for these works, there are other “benchmark-like” works on different aspects of this problem. For instance, Ganev et al. (Ganev et al., 2025) discuss the importance of discretization in tabular data synthesis. Moreover, Stadler et al. (Stadler et al., 2022) focus more on the quantitative evaluation of privacy gain. These works provide in-depth research from different perspectives but lack a more comprehensive viewpoint.
Realizing the weaknesses of current research, our work aims to address these issues by proposing a standardized algorithmic framework (Section 4), providing rigorous analysis (Section 5, Section 6 and Section 7), and conducting detailed experiments (Section 9).
@@ 4. Framework Overview
Some previous works (Liu et al., 2021; McKenna et al., 2021; Hu et al., 2023) have identified that current methods have several common working patterns and proposed unifying algorithmic frameworks, focusing primarily on feature selection and dataset synthesis. However, when dealing with a complex dataset, preprocessing it into a more manageable form should also be a necessary part of the algorithm. As presented Figure 1, our framework extends these approaches by incorporating a data preprocessing module. This section gives an overview of our framework and presents how modules work together to form a complete and cohesive data synthesis pipeline.
Preprocessing. As observed in our evaluation, many datasets contain attributes with large domain sizes (e.g., exceeding $10^{5}$ in Loan dataset (loa, 2020)), which pose significant challenges for data synthesis. For statistical methods, large attribute domains lead to expansive, low-dimensional marginals, which will introduce excessive DP noise during synthesis and degrade performance. For machine learning-based methods, a large domain requires a synthesizer with a substantial model size to learn the relationship between different attributes, leading to slower training and higher resource demands. Additionally, larger model sizes require a higher scale of DP noise under the same privacy budget compared to smaller synthesizers, degrading synthetic performance (Gong et al., 2025b).
Therefore, preprocessing is important and non-negligible in algorithm workflows, which can reduce the dimensionality of marginals by effectively merging those with similar characteristics and compressing the domains of the attributes. However, as shown in Table 1, preprocessing is often overlooked by prior works (Aydore et al., 2021; McKenna et al., 2022; Liu et al., 2021, 2023; Kotelnikov et al., 2023; Harder et al., 2021; Cai et al., 2021). This omission in the algorithm pipeline may affect the fairness of comparisons between methods.
Feature Selection. Even after effective preprocessing, the domain of the whole dataset increases exponentially with the number of attributes, making methods that rely on a histogram representation computationally infeasible. A practical approach is to utilize some representative local data features, such as low-dimensional marginals, to approximate the full joint data distribution (McKenna et al., 2022; Zhang et al., 2021; Cai et al., 2021; Liu et al., 2023). Therefore, the second step in the framework is to measure the representativeness of the features and select necessary features. To further analyze how to better select features in Section 6, we will categorize existing strategies into two categories:
@@
  2. Adaptive Feature Selection. Adaptive methods incorporate iterative calibration to refine the selection process. Typically, these methods begin with an initial feature selection, and then iteratively update the selected features based on intermediate feedback from previous selection steps.
@@
PrivSyn manually adjusts data records to match the marginals, while Private-GSD achieves this by genetic algorithm (Such et al., 2018). RAP and RAP++ use the relaxed projection mechanism to optimize the records of an initialized dataset. These methods are constructed by adjusting records. Some other methods fit models for data generation. PGM (McKenna et al., 2019) is a classical graphical model for tabular data synthesis, which utilizes a tree-like model to represent the distribution of data. Machine learning models such as generative networks and diffusion models have also been applied here.
@@ 5. Preprocessing
As mentioned in Section 3, when synthesizing complex datasets, we may encounter attributes with high cardinality. For example, an address or income attribute can have thousands of distinct values. Conducting statistics on such attributes is infeasible due to the high memory requirements and long execution times. Therefore, a preprocessing step is necessary for algorithm comparison. We surveyed some previous works (Zhang et al., 2021; McKenna et al., 2021; Ganev et al., 2025) and summarized the data types requiring preprocessing into two types: categorical and numerical. Before introducing this section, we need to clarify a basic assumption: domain information is considered public knowledge. This assumption is reasonable in many cases. For example, the domains of personal attributes in census data are well documented on the IPUMS website (IPUMS, 2025).
@@ 4       $b\leftarrow$ 1-way marginal of attribute $A_{j}$;
5       $\hat{b}=b+\mathcal{N}\left(0,\frac{1}{2\rho^{\prime}}\right)$;
6       for $i=1:|\hat{b}|$ do
@@ 5       $\hat{b}=b+\mathcal{N}\left(0,\frac{1}{2\rho^{\prime}}\right)$;
8                   replace $\hat{b}[i]$ with the rare encoding value;
@@ 5.1. Categorical Attributes Preprocessing
Some prior works (Zhang et al., 2021; McKenna et al., 2021) have proposed a $3\sigma$ merging strategy to preprocess categorical variables, where categories with counts below $3\sigma$ (where $\sigma$ represents the DP noise standard deviation in the counting process) are combined.
@@ 5.2. Numerical Attributes Preprocessing
Uniform Binning. Some previous works (Dick et al., 2023; Ganev et al., 2025) use uniform binning for continuous data preprocessing. It partitions an attribute’s domain into equal-length intervals, relying only on the attribute’s domain range and a predefined number of bins. Formally, this method can be expressed as Uniform $\text{Uniform Bin}(x)=\left\lfloor\frac{x-x_{\ell}}{h}\right\rfloor$, where $x_{\ell}$ is the lower bound of the attribute’s domain, and $h$ is the length of the uniform interval determined by the bin number.
@@ 8       Append $\mathcal{T}$ to $T$
PrivTree. A limitation of uniform binning is its reliance on a predetermined number of bins, which introduces concerns about hyperparameter selection. To address this, PrivTree decomposition (Zhang et al., 2016; Tao et al., 2022) can be used as a self-adaptive discretization method.
PrivTree employs a tree structure to iteratively divide the domain of an attribute, with splits continuing until intervals contain only a small number of records. However, since PrivTree utilizes sensitive data for domain division, it requires a fraction of the privacy budget to guarantee differential privacy. We use PrivTree on multiple attributes whose domain size is larger than a threshold, with algorithm details in Algorithm 2. We provide the proof of DP guarantee of this algorithm in the appendix.
@@ 6. Feature Selection
Selecting features determines the performance of many algorithms. Similar to preprocessing methods, it is important for statistical methods, while some machine learning methods can automatically learn the characteristics of the dataset by training models. In this section, we will briefly introduce the currently proposed selection methods and analyze them.
@@ 6.1. Existing Selection Methods
Some algorithms directly predefine a set of features instead of paying attention to selecting representative features. DP-MERF (Harder et al., 2021) leverages random Fourier features to capture correlations among numerical variables, while employing 2-way marginals for categorical variables. In addition, some methods with strong fitting ability can work on all two-way marginals. For example, Liu et al. (Liu et al., 2023) conduct experiments on all two-way marginals to demonstrate the performance of Private-GSD.
Instead of predefining some features, PrivSyn (Zhang et al., 2021) selects the most highly correlated marginals while respecting the privacy budget in one round. They measure each marginal using metric $\mathsf{InDif}_{i,j}=\left|M_{i,j}-M_{i}\times M_{j}\right|$, where $M$ denotes the attribute marginal. The selection process involves minimizing the expected error
@@
Adaptive Feature Selection. Different from non-adaptive methods, adaptive methods continuously update feature selection with feedback from the previous selection steps. In each synthesis round, adaptive feature selection conducts many queries on the current estimation, and the features with large errors are selected. These features will be used for the selection of future rounds. This strategy is used by methods like RAP (Aydore et al., 2021), RAP++ (Vietri et al., 2022), GEM (Liu et al., 2021), PrivMRF (Cai et al., 2021) and AIM (McKenna et al., 2022). These methods differ in several key aspects:
@@
  • Finally, the selection mechanism varies: AIM and GEM utilize the exponential mechanism (Cesar and Rogers, 2021), whereas RAP and RAP++ employ the Gumbel mechanism (Aydore et al., 2021), allowing them to select more than one feature in each round; PrivMRF takes a different approach, directly adding Gaussian noise to the selection criteria for selection and select the largest one.
Compared to the non-adaptive methods, the adaptive methods require multiple rounds of data synthesis or computation during the feature selection. Therefore, the efficiency of non-adaptive methods heavily depends on the efficiency of the data synthesis algorithms. Then, a natural question arises: which selection methods are theoretically better in utility? Or do the non-adaptive methods, which come at the cost of higher computational complexity, ensure continuous optimization in the correct direction during feature selection? We will discuss this in detail in the next section.
@@ 6.2. Analysis for Selection Algorithms
Some previous studies (McKenna et al., 2022; Cai et al., 2021; Liu et al., 2021) have discussed the important properties of a good selection mechanism. In this subsection, we formally investigate some aspects of it. Since all methods, except for DP-MERF, which predefines features without selection, focus on marginal selection, we only discuss marginal selection here.
@@
[…]
Therefore, we can compare these two errors, denoted as $\lVert M-M_{0}\rVert_{1}-n\sigma\sqrt{\frac{2}{\pi}}$, to determine the selection result, which demonstrates the importance of scale penalty term. This equation is also how some selection criteria involve the scale penalty term.
Superiority of Adaptive Selection. Without loss of generality, we consider the case of selecting 2-way marginals. Assuming that before selecting $(A_{i},A_{j})$, we have already fitted marginals $(A_{i},A_{1},\cdots,A_{k})$ and $(A_{j},A_{1},\cdots,A_{k})$, so that we can use this intermediate knowledge to estimate the distribution of $(A_{i},A_{j})$ as
@@
Before selecting the marginal $(A_{i},A_{j})$, the non-adaptive methods do not use intermediate results and can only use independent 1-way marginals to measure its representativeness. This independent measurement can be written as $\mathbb{D}_{\text{KL}}\left(\Pr[A_{i},A_{j}]\left\|\;\Pr[A_{i}]\Pr[A_{j}]\right.\right)$, where $\mathbb{D}_{\text{KL}}$ is the KL Divergence (Kullback and Leibler, 1951). The adaptive methods, because they select features based on intermediate synthesis results, will have a conditional estimation as $\mathbb{D}_{\text{KL}}\left(\Pr[A_{i},A_{j}]\left\|\;\hat{\Pr}[A_{i},A_{j}]\right.\right)$, which is the true KL divergence when choosing marginal $(A_{i},A_{j})$. Here $\hat{\Pr}[A_{i},A_{j}]$ is defined in Equation (1). Now we give the following theorem to prove the superiority of adaptive selection.
@@
$\mathbb{D}_{\text{KL}}\left(\Pr[A_{i},A_{j}]\left\|\;\hat{\Pr}[A_{i},A_{j}]\right.\right)\leq\mathbb{D}_{\text{KL}}\left(\Pr[A_{i},A_{j}]\left\|\;\Pr[A_{i}]\Pr[A_{j}]\right.\right)$
@@
The proof of Theorem 1 is deferred to the appendix. In essence, this theorem demonstrates that non-adaptive methods tend to overestimate the representativeness of features under KL divergence, whereas adaptive methods can correct this error by leveraging intermediate results.
@@ 7. Data Synthesis Module Comparison
GUM. GUM, used by PrivSyn, is an iterative adjustment method that modifies values in the initial dataset to align with the selected marginals. We assume that GUM merges marginals to $k$ cliques of sizes $\{c_{1},\cdots,c_{k}\}$. Since the maximum number of operations to fit each value in the marginals is no more than the size of the synthetic dataset, the time complexity of GUM is $\mathcal{O}\left(\sum_{i=1}^{k}Tc_{i}n\right)$, where $T$ is the number of update iterations, $n$ represents the synthetic dataset size, respectively. Because GUM strictly controls the clique size $c_{i}$, ensuring we have small cliques and simplifying the update process (e.g., we can choose a small $T$ to reach convergence). This guarantees the efficiency of GUM. However, it fits marginals one by one, overlooking overall correlations within the dataset, which may limit its utility.
Genetic Algorithm. Genetic algorithms (Such et al., 2018) use mutation and crossover operations to adjust datasets. The complexity depends on the number of mutations and crossovers per iteration. For instance, the algorithm by Liu et al. (Liu et al., 2023) has a complexity of $\mathcal{O}\left(T(P_{m}+P_{c})\right)$, where $T$ is the number of iterations, and $P_{m}$ and $P_{c}$ represent mutations and crossovers per iteration, respectively. The execution time of genetic algorithms is strongly influenced by the number of tuning rounds, which often exceeds the dataset size when handling complex datasets with diverse values.
Relaxed Projection. Relaxed projection mechanism (Aydore et al., 2021; Vietri et al., 2022) treats the dataset as a trainable model, optimizing it to match marginals. Given $T$ optimization rounds, synthetic data size $n$, and data dimension $d$, the time complexity is $\mathcal{O}(Tnd)$. This method can be regarded as both a model-fitting and data-adjusting approach. The complexity is driven by the size of the synthetic data and the number of optimization rounds. High-dimensional datasets with large attribute domains can result in an inflated encoded data dimension, making optimization difficult to converge and more time-intensive.
PGM. PGM (McKenna et al., 2019) constructs a junction tree of marginal cliques $C_{1},C_{2},\cdots,C_{k}$, ensuring that the intersection set $S_{i}$ of any two cliques appears only in those marginals. The overall distribution is approximated as:
@@
Generative Network. Generative networks (Liu et al., 2021) rely on parameters $m$, batch size $b$, and training iterations $T$. The training complexity is $\mathcal{O}(Tmb)$. Generating $n$ synthetic records results in a total complexity of $\mathcal{O}(Tmb+mn)$. The efficiency is affected by network design and training hyperparameters, which can be challenging for high-dimensional datasets.

### v2

@@ 3.2. Existing Benchmark Works
Realizing the weaknesses of current research, our work aims to address these issues by proposing a standardized algorithmic framework (Section 4), providing rigorous analysis (Section 5, Section 6, and Section 7), and conducting detailed experiments (Section 9).
@@ 4. Framework Overview
Some previous works (liu2021iterative; mckenna2021winning; hu2023sokprivacypreservingdatasynthesis) have identified that current methods have several common working patterns and proposed unifying algorithmic frameworks, focusing primarily on feature selection and dataset synthesis. However, when dealing with a complex dataset, preprocessing it into a more manageable form should also be a necessary part of the algorithm. As presented Figure 1, our framework extends these approaches by incorporating a data preprocessing module. This section gives an overview of our framework and presents how modules work together to form a complete and cohesive data synthesis pipeline.
Preprocessing. As observed in our evaluation, many datasets contain attributes with large domain sizes (e.g., exceeding $10^{5}$ in Loan dataset (loandata)), which pose significant challenges for data synthesis. For statistical methods, large attribute domains lead to expansive, low-dimensional marginals, which will introduce excessive DP noise during synthesis and degrade performance. For deep learning methods, a large domain requires a synthesizer with a substantial model size to learn the relationship between different attributes, leading to slower training and higher resource demands. Additionally, larger model sizes require a higher scale of DP noise under the same privacy budget compared to smaller synthesizers, degrading synthetic performance (gong2025dpimagebenchunifiedbenchmarkdifferentially).
Therefore, preprocessing is important and non-negligible in algorithm workflows, which can reduce the dimensionality of marginals by effectively merging those with similar characteristics and compressing the domains of the attributes. However, as shown in Table 1, preprocessing is often overlooked by prior works (aydore2021differentially; mckenna2022aim; liu2021iterative; liu2023generating; kotelnikov2023tabddpm; harder2021dp; cai2021data). This omission in the algorithm pipeline may affect the fairness of comparisons between methods.
Feature Selection. Even after effective preprocessing, the domain of the whole dataset increases exponentially with the number of attributes, making methods that rely on a histogram representation computationally infeasible. A practical approach is to utilize some representative local data features, such as low-dimensional marginals, to approximate the full joint data distribution (mckenna2022aim; zhang2021privsyn; cai2021data; liu2023generating). Therefore, the second step in the framework is to measure the representativeness of the features and select necessary features. To further analyze how to better select features in Section 6, we will categorize existing strategies into two categories:
@@
  2. Adaptive Feature Selection. Adaptive methods iteratively obtain new features based on intermediate feedback from previous selection steps to optimize algorithms. For example, AIM selects one new marginal at each iteration step and fits the model based on all selected marginals, which then serves as the reference for the next round of marginal selection. Moreover, DP-SGD iteratively calculates DP gradient features to update the model.
@@
PrivSyn manually adjusts data records to match the marginals, while Private-GSD achieves this by genetic algorithm (such2018deepneuroevolutiongeneticalgorithms). RAP and RAP++ use the relaxed projection mechanism to optimize the records of an initialized dataset. These methods are constructed by adjusting records. Some other methods fit models for data generation. PGM (mckenna2019graphical) is a classical graphical model for tabular data synthesis, which utilizes a tree-like model to represent the distribution of data. deep learning models such as generative networks and diffusion models have also been applied here.
@@ 5. Preprocessing
As mentioned in Section 3, when synthesizing complex datasets, we may encounter attributes with high cardinality. For example, an address or income attribute can have thousands of distinct values. Conducting statistics on such attributes is infeasible due to the high memory requirements and long execution times. Therefore, a preprocessing step is necessary for algorithm comparison. We surveyed some previous works (zhang2021privsyn; mckenna2021winning; ganev2025importancediscretemeasuringimpact) and summarized the data types requiring preprocessing into two types: categorical and numerical. Before introducing this section, we need to clarify a basic assumption: domain information is considered public knowledge. This assumption is reasonable in many cases. For example, the domains of personal attributes in census data are well documented on the IPUMS website (ipums).
@@ 5 $b\leftarrow$ 1-way marginal of attribute $A_{j}$;
6 $\hat{b}=b+\mathcal{N}\left(0,\sigma^{2}\right)$;
@@
    $\theta^{\prime}\leftarrow\max\left\{\theta\cdot\sum\hat{b},\;3\sigma\right\}$ ;
@@
9         replace $i$-th value with the rare encoding value;
@@ 5.1. Categorical Attributes Preprocessing
Some prior works (zhang2021privsyn; mckenna2021winning) have proposed a $3\sigma$ merging strategy to preprocess categorical variables, where categories with counts below $3\sigma$ (where $\sigma$ represents the DP noise standard deviation in the counting process) are combined.
@@ 5.2. Numerical Attributes Preprocessing
Uniform Binning. Some previous works (Dick_2023; ganev2025importancediscretemeasuringimpact) use uniform binning for continuous data preprocessing. It partitions an attribute’s domain into equal-length intervals, relying only on the attribute’s domain range and a predefined number of bins. Formally, this method can be expressed as $\text{Uniform Bin}(x)=\left\lfloor\frac{x-x_{\ell}}{h}\right\rfloor$, where $x_{\ell}$ is the lower bound of the attribute’s domain, and $h$ is the length of the uniform interval determined by the bin number.
@@ 8   Append $\mathcal{T}$ to $T$
PrivTree. A limitation of uniform binning is its reliance on a predetermined number of bins, which introduces concerns about hyperparameter selection. To address this, PrivTree decomposition (zhang2016privtree; tao2022benchmarkingdifferentiallyprivatesynthetic) can be used as a self-adaptive discretization method.
PrivTree employs a tree structure to iteratively divide the domain of an attribute, with splits continuing until intervals contain only a small number of records. However, since PrivTree utilizes sensitive data for domain division, it requires a fraction of the privacy budget to guarantee differential privacy. We use PrivTree on multiple attributes whose domain size is larger than a threshold, with algorithm details in Algorithm 2. We provide the proof of DP guarantee of this algorithm in the appendix of our full paper (chen2025benchmarking).
@@ 6. Feature Selection
Selecting features determines the performance of many algorithms. Similar to preprocessing methods, it is important for statistical methods, while some deep learning methods can automatically learn the characteristics of the dataset by training models. In this section, we will briefly introduce the currently proposed selection methods and analyze them.
@@ 6.1. Existing Selection Methods
Some algorithms directly predefine a set of features instead of paying attention to selecting representative features. DP-MERF (harder2021dp) leverages random Fourier features to capture correlations among numerical variables, while employing 2-way marginals for categorical variables. In addition, some methods with strong fitting ability can work on all two-way marginals. For example, Liu et al. (liu2023generating) conduct experiments on all two-way marginals to demonstrate the performance of Private-GSD.
Instead of predefining some features, PrivSyn (zhang2021privsyn) selects the most highly correlated marginals while respecting the privacy budget in one round. They measure each marginal using metric $\mathsf{InDif}_{i,j}=\left|M_{i,j}-M_{i}\times M_{j}\right|$, where $M$ denotes the attribute marginal. The selection process involves minimizing the expected error
@@
Adaptive Feature Selection. Different from non-adaptive methods, adaptive methods continuously update feature selection with feedback from the previous selection steps. In each synthesis round, adaptive feature selection conducts many queries on the current estimation, and the features with large errors are selected. These features will be used for the selection of future rounds. This strategy is used by methods like RAP (aydore2021differentially), RAP++ (vietri2022private), GEM (liu2021iterative), PrivMRF (cai2021data) and AIM (mckenna2022aim). These methods differ in several key aspects:
@@
  • Finally, the selection mechanism varies: AIM and GEM utilize the exponential mechanism (cesar2021bounding), whereas RAP and RAP++ employ the Gumbel mechanism (aydore2021differentially), allowing them to select more than one feature in each round; PrivMRF takes a different approach, directly adding Gaussian noise to the selection criteria for selection and select the largest one.
@@ 6.2. Analysis for Selection Algorithms
Some previous studies (mckenna2022aim; cai2021data; liu2021iterative) have discussed the important properties of a good selection mechanism. In this subsection, we formally investigate some aspects of it. Since all methods, except for DP-MERF, which predefines features without selection, focus on marginal selection, we only discuss marginal selection here.
@@
{ $\left\{\begin{aligned} &\text{ Selection error: }\;\lVert M-\hat{M}\rVert_{1}=n\sigma\sqrt{2/\pi}\\ &\text{ Unselection error: }\;\lVert M-M_{0}\rVert_{1}.\end{aligned}\right$.
Therefore, we can compare these two errors, denoted as $\lVert M-M_{0}\rVert_{1}-n\sigma\sqrt{\frac{2}{\pi}}$, to determine the selection result, which demonstrates the importance of the scale penalty term. This equation is also how some selection criteria involve the scale penalty term.
Superiority of Adaptive Selection. Establishing the superiority of the adaptive selection strategy is challenging in the absence of prior information about the generation module. Our analysis, therefore, proceeds under a set of reasonable assumptions. First, we consider the case of selecting 2-way marginals. Second, we assume that conditional independence is used for marginal selection. Specifically, assuming that before selecting $(A_{i},A_{j})$, we have already fitted marginals $(A_{i},A_{1},\cdots,A_{k})$ and $(A_{j},A_{1},\cdots,A_{k})$, so that we can use them to estimate the distribution of $(A_{i},A_{j})$ as
@@
[…]
@@
[…]
@@
The proof of Theorem 1 is deferred to the appendix of our full paper (chen2025benchmarking). In essence, this theorem demonstrates that under our assumptions, non-adaptive methods tend to overestimate the representativeness of features under KL divergence, whereas adaptive methods can correct this error by leveraging intermediate results. The more general comparisons between adaptive and non-adaptive selection strategies will be conducted in our experiments. Another notable fact is that compared to the non-adaptive methods, the adaptive methods require multiple rounds of data synthesis or computation during the feature selection, which can lead to higher time consumption as a trade-off for their superiority in utility.
@@ 7. Data Synthesis Module Comparison
GUM. GUM, used by PrivSyn, is an iterative adjustment method that modifies values in the initial dataset to align with the selected marginals. We assume that GUM merges marginals to $k$ cliques of sizes $\{c_{1},\cdots,c_{k}\}$. Since the maximum number of operations to fit each value in the marginals is no more than the size of the synthetic dataset, the time complexity of GUM is $\mathcal{O}\left(\sum_{i=1}^{k}Tc_{i}n\right)$, where $T$ is the number of update iterations, $n$ represents the synthetic dataset size, respectively. Because GUM strictly controls the clique size $c_{i}$, ensuring we have small cliques and simplifying the update process (e.g., we can choose a small $T$ to reach convergence). This guarantees the efficiency of GUM. However, it fits marginals one by one, overlooking overall correlations, limiting its utility.
Genetic Algorithm. Genetic algorithms (such2018deepneuroevolutiongeneticalgorithms) use mutation and crossover operations to adjust datasets. The complexity depends on the number of mutations and crossovers per iteration. For instance, the algorithm by Liu et al. (liu2023generating) has a complexity of $\mathcal{O}\left(T(P_{m}+P_{c})\right)$, where $T$ is the number of iterations, and $P_{m}$ and $P_{c}$ represent mutations and crossovers per iteration, respectively. The execution time of genetic algorithms is strongly influenced by the number of tuning rounds, which often exceeds the dataset size when handling complex datasets with diverse values.
Relaxed Projection. Relaxed projection mechanism (aydore2021differentially; vietri2022private) treats the dataset as a trainable model, optimizing it to match marginals. Given $T$ optimization rounds, synthetic data size $n$, and data dimension $d$, the time complexity is $\mathcal{O}(Tnd)$. This method can be regarded as both a model-fitting and data-adjusting approach. The complexity is driven by the size of the synthetic data and the number of optimization rounds. High-dimensional datasets with large attribute domains can result in an inflated encoded data dimension, making optimization difficult to converge and more time-intensive.
PGM. PGM (mckenna2019graphical) constructs a junction tree of marginal cliques $C_{1},C_{2},\cdots,C_{k}$, ensuring that the intersection set $S_{i}$ of any two cliques appears only in those marginals. The overall distribution is approximated as:
@@
Generative Network. Generative networks (liu2021iterative) rely on parameters $m$, batch size $b$, and training iterations $T$. The training complexity is $\mathcal{O}(Tmb)$. Generating $n$ synthetic records results in a total complexity of $\mathcal{O}(Tmb+mn)$. The efficiency is affected by network design and training hyperparameters, which can be challenging for high-dimensional datasets.
@@
Table 2. Summary of investigated datasets. We report the number of records, the number of attributes, numerical attributes, categorical attributes, and minimum and maximum attribute domain size.
Name #Records #Attr #Num #Cat Min/Max Domain
[…]